OpenStudy (migitmack):

Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? 3 7 17 21 Is she 21

OpenStudy (anonymous):

the answer is 17

OpenStudy (anonymous):

present age of mary=x,jane=y x=y-14; y+4=3(x+4) solving we get,x=3,y=17

OpenStudy (anonymous):

after 4 more years jane will be 21,so wait for 4 more years..;)